OpenStudy (anonymous):

rosa is 3 years older than her brother marco. let "x" stands for rosa's age and let "y" stands for marco's age . the linear equation of y= x -3 statse marcos age , y , in terms of rosas age , x . a.when marco is 38, write the equation for rosa's age. b.when marco is 72, write the equation for rosas age i need help!

OpenStudy (anonymous):

the original equation: y=x-3 If y is marco's age and that is 38 38=x-3, then just solve for x Again, if y is marco's age and that is 72 72=x-3, just solve for x

OpenStudy (anonymous):

so 73 -3 ?? Cx

OpenStudy (anonymous):

plug in marco's age for the y, and solve.

OpenStudy (anonymous):

can someone solve it for me TT-TT

OpenStudy (anonymous):

What you want is Rosa's age. The problem says x is Rosa's age. What you need to do is leave the x along on one side of the equation. From: 38=x-3 Since you have negative 3 along with the x, just add +3 to eliminate it (because -3+3=0) BUT, in order for the equation to remain valid, add +3 to the other side of the equation too. So: 38+3=x-3+3 And you have: 41=x And solve the other one in the same way.

OpenStudy (anonymous):

ok